# Explainable Procedural Mistake Detection

### v1

@@ 4.2.1 Coherence-Based Question Selection
$\texttt{Rel}(Q|T,\mathcal{Q},\mathcal{A})\max\limits_{A\in\{\text{{Yes}},\text{{No}}\}}\texttt{Inf}(A|Q,T,\mathcal{Q},\mathcal{A})$
@@
L 62.3 17.5 27.0 4.82 0.284
L 57.7 19.6 15.8 1.96 0.194
C 63.7 27.1 32.0 5.81 0.325
C 62.7 52.4 45.1 3.72 0.418
Table 1: Ego4D-PMD test set results for likelihood-based (L) and coherence-based (C) candidate question ranking approaches, with optional supplementary candidates generated through in-context learning (ICL).
@@
4.2.2 In-Context Learning Augmentation
Applying in-context learning in explainable PMD is not straightforward, as it would require reasoning over multiple images and possibly long, variable-length dialogs. Instead, we use in-context learning to improve the text-based VQG step by providing examples of human-written questions. We achieve this by annotating 20 procedures from the Ego4D training data with 3 reasonable questions one could ask about a given procedure to judge its success. [^9] These human-written questions achieve a mean 53.9% relevance and 83.2% maximum informativeness (i.e., for a yes or no answer). Procedures and questions in Appendix E.1. We prompt the VLM with these example procedures and questions, the current procedure at hand, and the previous 2 questions proposed by the VLM (as available) to incorporate information the VLM already collected. We then generate 4 additional candidate questions using the same constraints as described earlier. To minimize the impact of ordering, in-context examples are randomly shuffled in every prompt; a random seed is used to ensure examples are equivalently shuffled in different experiments.
4.2.3 Comparison of Results
We compare these approaches to vanilla VLMs based on mistake detection accuracy, mean example relevance and reference-adjusted informativeness of generated explanations (as defined in Section 3). We add two metrics for the efficiency of the self-dialog: the average number of iterations that occurred before stopping (i.e., the length of the dialog), and the average information gain in the success likelihood across all iterations in bits (i.e., how much information the VLM got from it).
@@ 4.2.3 Comparison of Results
Experimental results are presented in Table 2.101010Selected hyperparameters and validation set results for this table are listed in Appendix E.2. We find that coherence-based re-ranking and in-context learning [^11] To shed more light on where selected candidate questions come from in each approach, we visualize the distribution of question sources in Appendix E.3. both sharply improve the relevance and informativeness in all models, reaching a respective 74.2% and 46.0% in LLaVA. This suggests that questions that VLMs judge to be most likely are not naturally the most coherent. Interestingly, accuracy also jumps sharply for InstructBLIP and LLaVA to a maximum of 68.1%, demonstrating that coherent explanations are valuable to accurate PMD. Lastly, we find that the information gain under our proposed question selection interventions is consistently higher, reaching a maximum of 0.674 bits in LLaVA. This suggests that VLMs can make more confident conclusions given more coherent explanations.
4.3 Coherent Question Generation
While the above approaches boost the coherence of VLM outputs for PMD, they have limitations. First, since they are training-free, they do not improve the internal coherence of VLMs, rather they only filter and augment their outputs. Further, these interventions take significantly more time and compute due to the need to evaluate the coherence of candidate questions (which requires prompting an NLI model) and generate questions twice (once based on the self-dialog history and once with in-context learning). In practical applications like task guidance, it may be advantageous to use specialized VLMs to improve reliability and speed. As such, we lastly explore whether VLMs can be fine-tuned to generate more coherent questions using our automated coherence metrics.
Specifically, we apply direct preference optimization (DPO; Rafailov et al., 2023) with quantized low-rank adaptation (QLoRA; Hu et al., 2022; Dettmers et al., 2024) to fine-tune a VQG adapter for LLaVA (the best performing model). We generate training data for preference optimization by first running inference over the sampled Ego4D-PMD training data with coherence-based candidate question re-ranking and additional candidates generated through in-context learning. [^12] Here, we use a default set of inference hyperparameters $n=10$, $\delta=0.1$, and $\epsilon=0.05$. An ablation without in-context learning is presented in Appendix E.4. We then fine-tune VLMs on pairs of chosen and rejected candidate questions for each self-dialog turn based on their coherence-based ranking. [^13] Additional details, selected hyperparameters, and validation set results are listed in Appendix E.5. While the top ranked question is always used as the chosen question, the rejected question is randomly sampled from the bottom half of ranked questions. During explainable PMD evaluation, the trained adapter is applied to the VLM during VQG steps, and disabled for other steps.
4.3.1 Comparison of Results
@@
Table 2 displays the results of fine-tuning LLaVA for VQG, including the previous inference-time interventions to compare their relative impact on fine-tuned models. First, we observe that the relevance drastically increases from a maximum of 74.2% to 97.0%. Remarkably, this shows that VLMs are capable of learning to ask more coherent questions for explainable PMD, a task based on properties of the physical world.
@@
However, we find that accuracy and informativeness drop slightly from the best results in Table 2, suggesting that asking more coherent questions is not enough to improve performance globally. We suspect that this is due to an implicit difficulty trade-off of asking highly relevant questions, which should strongly indicate success if answered one way, otherwise a mistake. For elaborate procedures, highly relevant questions may cover multiple aspects or states of the scene. As such, answering these questions may be more difficult for VLMs, which struggle with complex questions. For example, while the generated question “Is the soil around the tomato seedling with the gardening trowel in your hand?” covers the success conditions of the procedure “Put some soil around the tomato seedling with the gardening trowel in your hand,” answering such a question requires understanding spatial relations between several objects (i.e., soil, tomato seedling, trowel, and hand). Future work may explore methods to encourage VLMs to generate simpler questions.
@@
Meanwhile, we find that fine-tuning reduces the average number of iterations taken to come to a conclusion by about 1, reaching a minimum of 1.72. Further, the average information gain reaches a new peak of 0.822. This suggests that coherence-based fine-tuning can empower VLMs to make more confident decisions faster. We lastly observe that the improvement in all metrics from coherence-based ranking and in-context learning is generally smaller in fine-tuned LLaVA than the base model (presented in Table 2). This suggests that our coherence metrics can improve the effectiveness of base VLMs for explainable PMD.  [^14] For a reference point with proprietary LMs, see Appendix E.6 for an evaluation of GPT-4o OpenAI et al. (2024).
@@ 4.4 Visualizing PMD Performance
An additional benefit of our automated coherence metrics is the ability to audit the global and local reasoning behaviors of VLMs under various strategies. In Figure 4, we visualize the distribution of decision error, relevance, and informativeness of four representative approaches to apply LLaVA. Specifically, we compare vanilla LLaVA (i.e., with likelihood-based question ranking) to variants successively equipped with coherence-based question ranking, in-context learning, and coherence-based fine-tuning with DPO.
For each example, decision error is calculated by how far the VLM’s success likelihood was from being 100% confident in the ground truth PMD decision. Relevance and informativeness use the example-level forms defined in Section 3. The colors of points indicate combinations of decision error, relevance, and informativeness in VLM outputs, highlighting common outcomes; we describe several examples in Appendix E.7. In comparing the plots, we see a dramatic shift as each successive intervention is introduced. Specifically, we see that red and black/indigo points are virtually eradicated, indicating fewer incoherent explanations and complete failures.
Vanilla LLaVA
+ Coherence Ranking
+ ICL
+ DPO
Ultimately, when choosing an approach to use for explainable PMD, one should weigh accuracy, coherence, and explanatory dialog efficiency as appropriate for the setting. For example, online applications may benefit from shorter (i.e., faster) dialog, even at a cost of accuracy. Meanwhile, high-risk applications may prioritize high confidence (e.g., information gain) over other factors. Additional insights toward the fine-grained strengths and weaknesses of various approaches may be gained from analyzing these results by mistake type, or verb and noun categories.
@@ 5 Related Work
In Section 1, we contextualized this work with prior work in PMD. In this section, we additionally discuss two lines of work: multi-step reasoning in LMs and leveraging NLI models in NLP.
5.1 Multi-Step Reasoning in LMs
Foundational LMs have exhibited impressive capabilities in multi-step reasoning from prompting-based methods Wei et al. (2022); Kojima et al. (2022). Later work strengthened LMs’ reasoning with multiple paths Wang et al. (2023b); Snell et al. (2024), tree-search Yao et al. (2023); Hao et al. (2023); Putta et al. (2024); Tian et al. (2024); Chen et al. (2024a); Zhang et al. (2024a); Qi et al. (2024), and fine-tuning on data generated by stronger LMs Wang et al. (2024); Gou et al. (2024). Related to task guidance, some work has investigated LMs’ ability to reason about dependencies between steps of physical procedures Bellos et al. (2024); Lal et al. (2024). In light of new challenges in visually-dependent reasoning Dai et al. (2023b); Li et al. (2023b); Guan et al. (2024); Zhang et al. (2024b), recent work has strengthened VLMs by utilizing text-based LMs and other tools to generate intermediate questions and coordinate visual reasoning step by step Srinivasan et al. (2024); Chen et al. (2024b); Zhou et al. (2024); Zong et al. (2024); Cheng et al. (2024). In this work, we similarly enabled VLMs to coordinate their visual reasoning through a self-dialog in the novel problem of explainable PMD, and we additionally proposed new coherence metrics to evaluate and improve their reasoning.
@@ 5.2 Leveraging NLI in Other NLP Tasks
NLI requires determining whether a premise text entails a hypothesis text, and has long been studied in the NLP community Dagan et al. (2005). Once thought to be a grand challenge for commonsense reasoning, many human-annotated resources have been compiled for NLI, and thus significant progress has occurred in training specialized LMs for NLI Storks et al. (2019). Prior work has leveraged this progress by using such NLI models to improve the competence, confidence, and coherence of LMs in other tasks, such as dialog systems Dziri et al. (2019); Welleck et al. (2019), summarization Roit et al. (2023), visual question answering Srinivasan et al. (2024), and image captioning Cascante-Bonilla et al. (2024). Our work extends this effort to a novel problem of explainable PMD, adopting two NLI-based coherence metrics.
@@ 6 Conclusion
In this work, we investigated the capabilities of foundational VLMs in a challenging new problem of explainable procedural mistake detection (PMD), where visual questions and answers must be generated to drive system decisions. To evaluate explanations, we defined two coherence metrics leveraging a fine-tuned NLI model. We then applied several interventions in VLMs to encourage coherent question selection (e.g., coherence-based question re-ranking) and generation (e.g., preference optimization from pairs of coherent and incoherent questions). Our results showed that vanilla VLMs do not generate coherent explanations off-the-shelf, but these interventions vastly improve their coherence, with the former also improving their overall PMD accuracy, and the latter improving the efficiency of generating and extracting information from explanations. Further, we showed that patterns in accuracy and coherence metrics illuminate common failures, e.g., visual hallucination. This work lays a foundation for future work harnessing VLMs for task guidance and other multimodal reasoning applications.
@@ 7 Limitations
One limitation of our explainable PMD problem formulation is the requirement of generating several pieces of information autoregressively, which would take several seconds in practical settings. This is not ideal for a problem like interactive task guidance, where responsiveness and the ability to intervene quickly to correct mistakes are important. However, rather than being applied frame by frame (which would likely not be feasible), we expect this process to be applied once at the end of procedure execution to verify the state of the environment, e.g., when the user asks a task guidance system to inform them of the next step of a recipe. Based on the results of our study, one could explore streamlined and specialized approaches to apply VLMs to a stream of video frames in a live online setting. For example, in preliminary experiments, we tried to generate questions once with a VLM, then answer them over a series of video frames, but we found this approach limited by the inability to adapt questions to previously gathered information, and the challenge of aggregating noisy VLM responses across time. We leave further investigation of such approaches for future work.

### v2

@@ 4.1.1 Coherence-Based Question Selection
$\texttt{Rel}(Q|T,\mathcal{Q},\mathcal{A})\mathop{\times\>\>\text{max}\>\>\texttt{Inf}}_{A\in\{\text{{Yes}},\text{{No}}\}}(A|Q,T,\mathcal{Q},\mathcal{A})$
@@
4.1.2 In-Context Learning Augmentation
Applying in-context learning in coherent PMD is not straightforward, as it would require reasoning over multiple images and dialogs of varying length. Instead, we use in-context learning to improve the text-based VQG step by providing examples of human-written questions. We achieve this by annotating 20 procedures from the Ego4D training data with 3 reasonable questions one could ask about a given procedure to judge its success. [^9] All questions and more details in Appendix E.1. We prompt the VLM with these example procedures and questions, the current procedure at hand, and the previous 2 questions proposed by the VLM (as available) to incorporate information the VLM already collected. We then generate 4 additional candidate questions using the same constraints described earlier. To minimize the impact of ordering, in-context examples are randomly shuffled in every prompt. [^10] Examples are equivalently shuffled in compared results.
4.2 Coherent Question Generation
While the above approaches may boost the coherence of PMD rationales, they have limitations. First, since they are training-free, they do not improve the internal coherence of VLMs, rather they only filter and augment their outputs. Further, these interventions take significantly more time and compute due to the need to evaluate the coherence of candidate questions (which requires prompting an NLI model) and generate questions twice (once based on the self-dialog history and once with in-context learning). In practical applications like task guidance, it may be advantageous to use specialized VLMs to improve reliability and speed. As such, we also explore whether VLMs can be fine-tuned to generate more coherent questions using our automated coherence metrics.
Specifically, we apply direct preference optimization (DPO; Rafailov et al., 2023) with quantized low-rank adaptation (QLoRA; Hu et al., 2022; Dettmers et al., 2024) to fine-tune a VQG adapter for LLaVA. We generate training data by first running inference over the Ego4D-PMD training data with coherence-based candidate question re-ranking and additional candidates generated through in-context learning. [^11] Ablation without in-context learning in Appendix E.2. We then fine-tune VLMs on pairs of chosen and rejected candidate questions for each self-dialog turn based on their coherence ranking score. [^12] More details and validation results in Appendix E.3. The top ranked question is always chosen, while a rejected question is sampled from the bottom half of candidates. At inference time, the trained adapter is applied during VQG, and disabled for other steps.
5 Experimental Results
In this section, we evaluate VLMs on coherent PMD off-the-shelf and under the previously introduced interventions. Metrics include mistake detection accuracy, mean example relevance and reference-adjusted informativeness of rationales (defined in Section 3). We add two metrics for self-dialog efficiency: the average number of iterations that occurred before stopping (i.e., the length of the dialog), and the average information gain in the success likelihood across all iterations in bits (i.e., how much information the VLM got from it). We lastly show that our proposed metrics visualize common behaviors of VLMs, enabling a panoptic understanding of their performance.
We specifically evaluate InstructBLIP Dai et al. (2023a), LLaVA 1.5-7B Liu et al. (2023), and Llama 3.2-Vision-11B Dubey et al. (2024), small open-source VLMs feasible for online use (important for real-world applications like task guidance). They apply different architectures and training strategies to integrate vision into their LMs. LLaVA and InstructBLIP were not trained on Ego4D, while Llama 3’s training data was not disclosed. For a reference point with state-of-the-art proprietary VLMs, we additionally include results with off-the-shelf GPT-4o OpenAI et al. (2024b). Additional details and prompt templates can be found in Appendix D and E.4.131313In Appendix E.5, we include an additional evaluation of VLMs without generating rationales (as done in prior work).
@@ 5 Experimental Results
L 61.0 16.5 .275 4.70 .223
L 59.1 15.9 .317 6.51 .256
C 60.2 25.2 .341 6.35 .264
C 61.7 52.5 .436 3.59 .379
GPT-4o
L 55.4 54.0 .220 1.84 .793
Table 1: Ego4D-PMD test set results for GPT-4o OpenAI et al. (2024b) and likelihood-based (L) and coherence-based (C) candidate question ranking approaches, with optional supplementary candidates generated through in-context learning (ICL).
@@
5.1 Coherent Question Selection Results
@@
Experimental results for coherent question selection interventions are presented in Table 2.141414Hyperparameters and validation results in Appendix E.6. We find that coherence-based re-ranking and in-context learning [^15] We visualize the distribution of question sources (i.e., in-context learning vs. full self-dialog history) in Appendix E.7. both sharply improve the relevance and informativeness in all models, reaching a respective 75.5% and 0.464 bits in LLaVA. This suggests that questions that VLMs find most likely are not naturally the most coherent. Interestingly, accuracy also jumps sharply for InstructBLIP and LLaVA to a maximum of 67.8%, demonstrating that coherent rationales are valuable to accurate PMD. The information gain under our question selection interventions is consistently higher, reaching a maximum of 0.663 bits in LLaVA. This suggests that VLMs can make more confident conclusions given more coherent rationales. [^16] To better understand performance gains here, in Appendix E.8, we contextualize these results with a naïve semantic diversity-based ranking baseline. Meanwhile, our best configurations outperform GPT-4o in accuracy, relevance, and informativeness at less than 10% of its size Abacha et al. (2025). While GPT-4o takes fewer iterations and exhibits higher information gain (thus making faster and more confident decisions), its rigidity as a closed proprietary VLM prevents further improvements.
5.2 Coherent Question Generation Results
@@
When fine-tuned for coherent question generation, Table 2 shows that relevance drastically increases from a previous maximum of 75.5% to 95.0%. Remarkably, this demonstrates that VLMs can learn to ask more coherent questions for PMD, a task based on properties of the physical world.
@@
However, we find that accuracy and informativeness drop slightly from the best results in Table 2 to a respective maximum of 64.2% and 0.340 bits. This suggests that asking more relevant questions is not enough to improve performance globally. We suspect this is due to an implicit difficulty trade-off of asking highly relevant questions, which strongly indicate success if answered one way, otherwise a mistake. For elaborate procedures, highly relevant questions may cover multiple aspects or states of a scene. As such, answering these more complex questions may be difficult for VLMs. For example, while the VLM-generated question “Is the soil placed around the seedling with the trowel in the person’s hand?” covers the success conditions of the procedure “Put some soil around the tomato seedling with the gardening trowel in your hand,” answering it requires understanding spatial relations between several objects (i.e., soil, seedling, trowel, and hand). Future work may explore methods to encourage VLMs to generate simpler questions. [^17] As an initial inquiry here, we experiment with applying an exponential length penalty to VQG in Appendix E.9.
@@
Meanwhile, we find that fine-tuning reduces the average number of iterations taken to come to a conclusion by about 1.25, reaching a minimum of 1.81 (matching that of GPT-4o). This is especially important for online PMD, which requires a fast reaction. Further, the average information gain reaches a new peak of 0.742, nearly matching that of GPT-4o. This suggests that coherence-based fine-tuning can empower VLMs to make more confident decisions faster. We lastly observe that coherence-based ranking and in-context learning have a smaller impact on performance metrics in fine-tuned LLaVA than in the base model (Table 2). This suggests that fine-tuning VLMs using our coherence metrics reduces the need for inference-time interventions for coherence.
Vanilla LLaVA
+ Coherence-Based Ranking
+ In-Context Learning
+ Coherence-Based DPO
@@ 5.3 Visualizing PMD Performance
An additional benefit of our automated coherence metrics is the ability to audit the reasoning behaviors of VLMs. In Figure 4, we visualize the distribution of decision error, example-level relevance, and example-level informativeness of four representative approaches. Specifically, we compare vanilla LLaVA (i.e., with likelihood-based question ranking) to variants successively equipped with coherence-based question ranking, in-context learning, and coherence-based fine-tuning with DPO. For each example, decision error is calculated by how far the VLM’s success likelihood is from being 100% confident in the ground truth success decision. Point colors indicate combinations of decision error, relevance, and informativeness, highlighting common local outcomes. Several examples are described in Appendix E.10.
In comparing the plots globally, we see a virtual elimination of red, black, and indigo points with irrelevant and uninformative rationales. We also see a significant reduction in the range of informativeness under coherence-based fine-tuning. This suggests that when questions are answered correctly under this approach, it is highly informative to the success decision, and vice versa. This is in line with our earlier observation that coherence-based fine-tuning may encourage the generation of overly relevant and thus complex questions, which can make or break the rationale. There are also a large number of points with zero informativeness, i.e., unsure answers to these complex questions.
@@ 6 Related Work
Beyond the prior work in PMD discussed in Section 1, we next discuss other related research.
6.1 Multi-Step Reasoning in LMs & VLMs
LMs have exhibited impressive reasoning capabilities from prompting methods Wei et al. (2022); Kojima et al. (2022), later strengthened with multiple paths Wang et al. (2023b); Snell et al. (2024), tree-search Yao et al. (2023); Hao et al. (2023); Putta et al. (2024); Tian et al. (2024); Chen et al. (2024a); Zhang et al. (2024); Qi et al. (2024), and fine-tuning on reasoning chains from stronger LMs Wang et al. (2024); Gou et al. (2024); Muennighoff et al. (2025) and/or self-generated during reinforcement learning OpenAI et al. (2024a); DeepSeek-AI et al. (2025). Related to PMD, some work has investigated LMs’ reasoning about dependencies between physical procedures Bellos et al. (2024); Lal et al. (2024). Meanwhile, other work has applied iterative self-questioning approaches to deepen inquiry in other domains, e.g., medicine and fact verification Cattan et al. (2024); Li et al. (2025); Vladika et al. (2025).
In light of challenges in visual reasoning in VLMs Dai et al. (2023b); Li et al. (2023b); Guan et al. (2024), some work has proposed training-free strategies and training paradigms to reduce visual hallucination Wan et al. (2024); Leng et al. (2024); An et al. (2024); Ganz et al. (2024), and utilized LMs and other tools to generate intermediate questions and coordinate visual reasoning step by step You et al. (2023); Srinivasan et al. (2024); Chen et al. (2024b); Zhou et al. (2024); Zong et al. (2024); Cheng et al. (2024); Zhu et al. (2024); Sun et al. (2024); Jing and Rahman (2025). The latter stems from prior efforts in visual dialog Das et al. (2017); Kottur et al. (2019), which has been used to study and improve machine performance in various problem areas Chattopadhyay et al. (2017); de Vries et al. (2017); Uehara et al. (2022); Wang et al. (2022). Other works have assessed (V)LMs’ reasoning about procedures based on visual information Yang et al. (2018); Hendricks and Nematzadeh (2021); Jin et al. (2022); Yuksekgonul et al. (2023); Nguyen et al. (2024, 2025). In this work, we evaluated self-generated visual dialogs from VLMs in a novel problem of coherent PMD.
@@ 6.2 Leveraging NLI in Other NLP Tasks
NLI requires determining whether a premise text entails a hypothesis text, a reasoning challenge long studied in NLP Dagan et al. (2005). Many human-annotated resources have been compiled for NLI, thus significant progress has occurred in developing NLI models Storks et al. (2019). Prior work has leveraged this progress by using NLI models to improve the competence, confidence, and coherence of LMs in other tasks, such as dialog systems Dziri et al. (2019); Welleck et al. (2019), summarization Roit et al. (2023), VQA Srinivasan et al. (2024), and image captioning Cascante-Bonilla et al. (2024). Building on this, we adopt two NLI-based metrics to strengthen VLMs in the novel problem of coherent PMD.
@@ 7 Conclusion
In this work, we evaluated foundational VLMs on a challenging problem of coherent PMD, where visual questions and answers must be generated to drive success decisions. To evaluate these rationales, we leveraged an NLI model to define two coherence metrics, using them to encourage coherent question selection and generation through common interventions. Our results showed that VLMs do not generate coherent rationales off-the-shelf, but these interventions improve their coherence, with the former also improving accuracy, and the latter improving the efficiency of generating and extracting information from rationales, albeit creating tradeoffs between these aspects. Further, patterns in accuracy and coherence metrics revealed detailed performance insights. Ultimately, when choosing an approach for coherent PMD, accuracy, coherence, and efficiency must be weighed for the setting. For example, while online applications may require shorter dialog (even at a cost of accuracy), high-risk applications may prioritize accuracy and confidence. This work lays a foundation for future research in the application of VLMs to PMD and task guidance, and investigation of how these system goals interact.
@@ 8 Limitations
One limitation of our coherent PMD problem formulation is the requirement of generating several pieces of information autoregressively, which would take several seconds in practical settings. This is not ideal for a problem like interactive task guidance, where responsiveness and the ability to intervene quickly to correct mistakes are important. However, rather than being applied frame by frame (which would likely not be feasible), we expect this process to be applied once at the end of procedure execution to verify the state of the environment, e.g., when the user asks a task guidance system to inform them of the next step of a recipe. Based on the results of our study, one could explore streamlined and specialized approaches to apply VLMs to a stream of video frames in a live online setting. For example, in preliminary experiments, we tried to generate questions once with a VLM, then answer them over a series of video frames, but we found this approach limited by the inability to adapt questions to previously gathered information, and the challenge of aggregating noisy VLM responses across time. We leave further investigation of such approaches for future work.
@@
Lack of ground truth rationales.
Lastly, it is worth noting that the dataset collected does not include ground truth rationales for mistake detection labels. Instead, we opted to propose automated coherence metrics for generated questions and answers based on a fine-tuned NLI model, which itself is prone to error and thus limits the objectivity of our evaluations. We chose this path for two important reasons. First, there may be multiple valid ways to detect a mistake through asking and answering visual questions, each of which could consist of different questions and/or different numbers of questions. [^18] For example, in trying to determine the success of the procedure “In a bowl, add the cut cherry tomatoes” from Peddi et al. (2024), we could reasonably ask one question “Are all the cherry tomatoes in the bowl?” or two questions “Are there cherry tomatoes in the bowl?” and “Are there any cherry tomatoes outside of the bowl?” In our opinion, existing metrics for text generation (which largely measure syntactic or semantic similarity of text) are not as well suited for this extremely challenging evaluation as fine-tuned NLI models (which are optimized to judge the logical consistency between texts). Second, in a real-world setting like PMD for task guidance, we believe that automated metrics are better suited for continually understanding and improving a deployed system than an offline benchmark of ground truth rationales. Despite this limitation, we believe that the value of these metrics is demonstrated by the accuracy and efficiency improvements brought by incorporating them into coherent PMD.